# Learning Tractable Distributions of Language Model Continuations

### v1

@@ 3.1 Inference using Hybrid HMMs
As a concrete application of the tractability of a neural-encoded HMM, we consider the task of generated text conditioned on a (logical or semantic) constraint $\alpha$. A natural decoding procedure is to sample autoregressively from the conditional distribution factored by Bayes’ rule as Equation 3, where the first term $p(x_{t}|x_{<t})$ is computed with the autoregressive model, and the second term $p(\alpha|x_{\leq t})$ is computed with the tractable neural-encoded HMM serving as a proxy for the autoregressive model. As an example, we describe the case where $\alpha$ is the event that the text is accepted by a given deterministic finite automaton (DFA); see Appendix B for the definition of a DFA. We follow closely the algorithm Ctrl-G designed for unconditional HMMs (ZhangNeurIPS24). Given a DFA $M$, let $S_{t}$ be the random variable representing the state of $M$ after reading $x_{\leq t}$ sampled from $p(x_{\leq t}|\alpha)$. Then,
$p(\alpha|x_{\leq t})$ $=\sum_{z_{t}}p(z_{t}|x_{\leq t})p(\alpha|z_{t},x_{\leq t})=\sum_{z_{t}}p(z_{t}|x_{\leq t})p(\alpha|z_{t},s_{t})$ (8)
where the first equality follows from the law of total probability, and the second from the Markov properties of HMMs and DFAs and the fact that $s_{t}$ is fully determined by $x_{\leq t}$. The term $p(z_{t}|x_{\leq t})$ may be estimated by the neural encoder. In practice, to avoid evaluating the neural encoder $V$ times, once for each candidate next token $x_{t}$, we instead evaluate the encoder once to compute $p(z_{t-1}|x_{<t})$ and then perform a single HMM forward step to obtain $p(z_{t}|x_{\leq t})$ for each $x_{t}$.
Then, observe that the term $p(\alpha|z_{t},s_{t})$ is entirely independent of the context $x_{\leq t}$, and so all $TVH$ such probabilities $p(\alpha|z_{t},s_{t})$ can be precomputed and stored in a lookup table (before decoding); there is an efficient backward-style algorithm for this computation (ZhangNeurIPS24). Finally $p(\alpha|x_{\leq t})$ is computed by the aggregation in Equation 8. Similar approaches work for other constraints. TRACE is a similar algorithm derived by YidouWengICML25 for the case where the constraint is an attribute $s$ predicted by a fully factorized probabilistic classifier $p(s|x_{1:n})$.
@@
We emphasize an advantage to changing only the encoder of the HMM. Specifically, we contrast neural-encoded HMMs with an alternative wherein a neural network predicts, given a context $x_{\leq t}$, fresh parameters of a full HMM. With such full conditioning, the backward computation of $p(\alpha|z_{t},s_{t})$ can no longer be precomputed but needs to be carried out at each step of decoding on the freshly predicted HMM, resulting in a decoding time that scales quadratically in sequence length. Moreover, fresh HMM parameters are needed for each context, resulting in an additional blowup in a memory with batch size. The time and space complexities for training and inference of standard HMMs, neural-encoded HMMs, and full conditional HMMs are summarized in Appendix C. We use the parameter $\tau$ for the complexity of a single HMM forward (or backward) step (always assuming $\tau=\Omega(hV)$); for standard (dense) transition and emission matrices, $\tau=O(h^{2}+hV)$.
@@ 3.2 Architecture
We now specify the architectural designs for our hybrid model, which includes two components: the HMM decoder and the neural network encoder.
@@ 3.2.1 Structured Sparsity
Proposition 1
For any Markov chain $X_{<t}\to Z_{t}\to X_{\geq t}$, we have
$I(X_{<t};X_{\geq t})\leq H(Z_{t})\leq\log|\text{supp}\,(Z_{t})|.$
This bound (proven in Appendix A) holds regardless of the encoder $p(z_{t}|x_{<t})$ or decoder $p(x_{\geq t}|z_{t})$, i.e., it holds for standard HMMs as well as neural-encoded HMMs. Although Proposition 2 provides a strong motivation for increasing the hidden size $H$ to improve the model’s capacity, the number of parameters in an HMM with dense transition and emission matrices grows quadratically in hidden size. Therefore, we also consider using Monarch matrices (dao2022monarch), structured matrices that require polynomially fewer parameters in the hidden size yet can still express complex, high-rank matrices, which have been used for scaling tractable models (ZhangICML25). Monarch matrices for the transition and emission matrices in an HMM use a combined $O(h^{3/2}+h^{1/2}V)$ set of parameters, giving a single-step complexity of $\tau=O(h^{3/2}+h^{1/2}V)$.
Figure 2: Neural architectures for neural-encoded HMMs: (a) frozen Transformer with linear mapping, (b) frozen Transformer with additional learnable layer, (c) fully finetuned Transformer.
@@ 3.2.2 Neural Network Architectures
Another key factor that determines the performance of neural-encoded HMMs is the choice of the neural network. Aside from achieving better conditional log-likelihoods, we want the neural encoder to be as lightweight as possible to minimize the computational overhead in downstream applications. One way to achieve this is by reusing the same neural network backbone of existing autoregressive sequence models (e.g., LLMs) and adding minimal modules on top of them.
Figure 2 illustrates the three neural architectures we adopted. The first one (in sub-figure (a)) keeps the (causal) Transformer backbone frozen and only adds a learnable linear mapping on top of its last hidden embedding to predict the latent variable distribution $p(z_{t}|x_{\leq t})$. In downstream applications, the Transformer embeddings are already computed, so we only add a negligible overhead of running the linear layer. The second variant (in sub-figure (b)) extends the frozen Transformer with an additional learnable Transformer layer before the output linear mapping, which strikes a better balance between the expressiveness of the encoder and the computational overhead. The third design (in sub-figure (c)) fully finetunes the Transformer backbone together with the linear prediction layer. While this variant has the largest number of trainable parameters, it also provides the most flexibility to tailor the representation for the HMM.
@@ 4 Experiments
Datasets
In our experiments, we utilize datasets consisting of data sampled from LLMs. Our main evaluation is on datasets from GPT2-large, GPT2-large finetuned for CommonGen, and Qwen2.5-VL-2B. For each of these datasets, we unconditionally sample sequences of length $T=32$, and uniformly at random choose a split position in the sequence to construct context and continuation pairs of variable length.
Methods and Baselines
In the distillation performance experiments, we primarily compare our neural HMMs to corresponding HMMs in terms of their performance in modeling continuations. For reference, we also compare to HMMs trained using the method of lee2025optimizing. In the controlled generation benchmarks, we compare LTLA with the standard HMMs used in Ctrl-G, in order to show the benefits of the neural encoder, and with prompting and sampling baselines for the novel VLM detoxification setting.
Research Questions
The key research questions we seek to investigate are as follows. (RQ1) Do neural HMMs achieve better modeling performance compared to standard HMMs, and if so, in what way? (RQ2) How does the choice of neural encoder architecture affect these results? (RQ3) Do any gains in terms of modeling performance translate to improvements in downstream benchmarks via LTLA? (RQ4) Can neural HMMs effectively condition on multimodal context data?
Figure 3: Perplexity of neural-encoded HMMs and baseline HMM for varying hidden sizes with dense transition and emission matrices on the left, Monarch matrices on the right.
@@ 4.1 Distillation Performance & Analysis
Figure 3 reports the scaling behavior of neural-encoded HMMs compared to standard HMM baselines. The a-axis refers to the hidden size of the HMM. Across both dense and Monarch parameterizations (as introduced in Section 3.2.1), augmenting HMMs with neural components consistently reduces perplexity relative to the baseline. Across the three neural architectures, adding a trainable Transformer layer (i.e., Figure 2(b)) or finetuning GPT (i.e., Figure 2(c)) achieves stronger performance compared to the learnable linear layer. These results demonstrate that neural HMMs scale favorably with hidden dimension, leading to better conditional log-likelihood per token/perplexity. In Figure 4, we further see that the neural encoder achieves better perplexity particularly for shorter continuation lengths, which is to be expected as the first few tokens have the strongest dependence on the context and are where the neural encoder has the most direct effect.
@@
For reference, we also compared to the method of lee2025optimizing, who conduct an empirical analysis of techniques designed to improve the learning dynamics of HMM language models, and suggest utilizing a neural reparameterization (chiu2020scaling) combined with latent variable distillation (LiuICLR23). Contrary to their findings, we found that this worsens performance over standard training in our distillation setting. In particular, we found that the performance of the HMM is quite sensitive to the hyperparameters of the AdamW optimizer, but by choosing appropriate hyperparameters it outperforms the neural parameterization and LVD.
@@ 4.2 Controlled Generation
LTLA delivers (i) improved distillation quality and (ii) compatibility with autoregressive models beyond the reach of a standalone TPM. The stronger predictive fit of our neural-encoded HMMs translates into better performance on standard LLM controlled-generation benchmarks. In addition, LTLA’s flexible design allows us to extend this tractable lookahead approach to VLMs. For VLM image captioning, LTLA uses its tractable lookahead to compute the probability of a future semantic property (i.e. toxicity) that is then used to reweight the VLM’s original next-token distribution. In our experiments, this approach outperforms sampling-based and heuristic controllers with negligible inference overhead. We demonstrate versatility across two comprehensive constraint families—hard logical constraints and soft semantic attributes—using the established Ctrl-G (ZhangNeurIPS24) and TRACE (YidouWengICML25) frameworks. By conditioning the TPM on transformer prefixes and learning to look ahead better, LTLA improves constraint satisfaction while maintaining the base model’s fluency and diversity.
@@
Logical Constraints We enforce hard constraints (keywords, propositions) by combining the tractable lookahead model with deterministic finite automata (DFAs) in the Ctrl-G framework (ZhangNeurIPS24). We distill both standard HMMs and LTLA variants from a GPT-2 model fine-tuned on CommonGen. For each keyword set, we construct a DFA that accepts only sequences satisfying the constraints, then run Ctrl-G beam search (128 beams) and select the top hypothesis by base-LM log-likelihood. We evaluate against references using BLEU-4, CIDEr, and ROUGE-L, and report average and maximum perplexity for fluency.
Table 2: Generation quality of HMM variants on the CommonGen benchmark using Ctrl-G.
Model BLEU-4 Cider RougeL Avg. Perplexity($\downarrow$) Max Perplexity ($\downarrow$)
Max Sequence Length 20
Standard HMM 0.301 1.552 0.448 39.59 1569.51
HMM with Linear NN 0.303 1.566 0.448 41.99 1458.38
HMM with Transformer Block 0.297 1.536 0.446 36.24 616.73
Max Sequence Length 32
Standard HMM 0.303 1.566 0.448 36.00 1569.51
HMM with Linear NN 0.311 1.566 0.448 34.16 671.88
HMM with Transformer Block 0.305 1.573 0.448 33.98 1065.47
As shown in Table 2, LTLA leads to significant improvements in generation quality under hard constraints. Neural HMMs not only demonstrate consistent syntactical gains over the standard HMM across metrics including BLEU-4, Cider, and RougeL, but also achieve significant fluency gains by dramatically reducing perplexity, slashing the maximum perplexity by more than half. This suggests that the neural encoders are highly effective at providing more context-aware guidance for the Ctrl-G framework to satisfy constraints without forcing the model to generate unnatural sequences.
Semantic constraints in VLMs. We introduce a plug-and-play tractable lookahead for VLMs, compatible with autoregressive decoders where a standalone TPM is ill-suited. LTLA enforces soft constraints such as style (emotion, sentiment) and topicality with only the addition of a log-linear classifier, making it extremely lightweight and adaptable for new constraints. We adopt TRACE (YidouWengICML25) to detoxify image captioning on the Hateful Memes dataset (kiela2020hateful).
@@
Concretely, we distill LTLA from Qwen2-VL-Instruct-2B and pair it with a log-linear toxicity model trained on RealToxicityPrompts (gehman2020realtoxicityprompts), scored with the Perspective API (lees2022new). The predicted future toxicity reweights the next-token distribution during decoding, providing probabilistic guidance without architectural changes.
Baselines. We compare against (i) default Qwen generation, (ii) prompt engineering with “Avoid any toxic, offensive, or hateful language.”, and (iii) Monte Carlo sampling that reweights tokens using the mean sampled toxicity of $N$ continuations per step.
Results. As shown in Table 3, LTLA achieves the strongest detoxification while preserving fluency. Under a comparable compute budget, the Monte Carlo (MC) baseline was capped at k=1k{=}1 reweighting pass per step, which still ran at approximately $10\times$ the LTLA decoding time. Both methods reweight continuations to approximate the same future-toxicity query, MC by sampling and LTLA via a tractable predictive surrogate, but at equal budget the MC estimator is high variance and costly, yielding higher toxicity and lower fluency..
Table 3: Toxicity vs. fluency across methods. Lower is better.
Method Avg. toxicity ($\downarrow$) Avg. max toxicity ($\downarrow$) Perplexity ($\downarrow$)
Vanilla 0.087 0.249 2.359
Neural Monarch HMM 0.064 0.188 3.564
Prompting 0.078 0.224 2.463
Sampling 0.097 0.309 43.142
@@ 5 Related Work and Conclusion
We study the problem of computing conditional queries on autoregressive sequence models (boyd2022predictive), which has many existing approaches, especially for constrained generation (qin2022cold; hu2023amortizing; lew2023sequential; krause2020gedi; yang2021fudge; meng2022controllable). We follow a recent line of work, proposing a method based around tractable modeling (ZhangICML23; ZhangNeurIPS24; YidouWengICML25), adopting hidden Markov models as our primary structure (rabiner2003introduction; chiu2020scaling; lee2025optimizing). Our amortized inference technique of neurally conditioning on context is similar to hybrid neural and tractable models in other settings (shao2020conditional; dos2024probabilistic). Our work complements existing research in tractable modeling, that aims to scale tractable models in practice (loconte2024subtractive; LiuICML24; maeneklay; wang2025relationship) while characterizing theoretical limits (ZhangICML21; blaser2023not; harviainen2023inference; BroadrickUAI24).
In conclusion, we proposed Learning to Look Ahead (LTLA), a novel approach to modeling and probabilistically reasoning about language model continuations using tractable models. LTLA uses a neural encoder to predict the latent state of a tractable hidden Markov model, enabling more accurate dependence on the context while maintaining the tractability and computational efficiency of the approach. Empirically, we show that LTLA significantly boosts the perplexity of tractable models over the prior state-of-the-art, particularly for shorter continuation sequences, and improved performance on controllable generation benchmarks.

### v2

@@ 3.1 Inference using Neural HMMs
As a concrete application of the tractability of a neural HMM, we consider the task of generating text conditioned on a (logical or semantic) constraint $\alpha$. We sample autoregressively from the conditional distribution factored as in Equation 2, where the term $p(x_{t}|x_{<t})$ is computed by the autoregressive model, and the term $p(\alpha|x_{\leq t})$ is computed with the tractable neural HMM. As an example, we describe the case where $\alpha$ is the event that the text is accepted by a given deterministic finite automaton (DFA); see Appendix B for a full definition. We follow the Ctrl-G algorithm designed for unconditional HMMs (Zhang et al., 2024). Given a DFA $M$, let $S_{t}$ be the random variable representing the state of $M$ after reading $x_{\leq t}$ sampled from $p(x_{\leq t}|\alpha)$. Then,
$p(\alpha|x_{\leq t})=\sum_{z_{t}}p(z_{t}|x_{\leq t})p(\alpha|z_{t},s_{t})$ (7)
using the law of total probability together with the Markov properties of HMMs and DFAs (and the fact that $s_{t}$ is fully determined by $x_{\leq t}$). The term $p(z_{t}|x_{\leq t})$ may be estimated by the neural encoder. In practice, to avoid evaluating the neural encoder $V$ times, once for each candidate next token $x_{t}$, we instead evaluate the encoder once to compute $p(z_{t-1}|x_{<t})$ and then perform a single HMM forward step to obtain $p(z_{t}|x_{\leq t})$ for each $x_{t}$. Moreover, observe that the term $p(\alpha|z_{t},s_{t})$ is entirely independent of the context $x_{\leq t}$, and so all $T\cdot V\cdot H$ such probabilities $p(\alpha|z_{t},s_{t})$ can be precomputed and stored in a lookup table before decoding via an efficient backward-style algorithm (Zhang et al., 2024). Similar approaches work for other constraints; TRACE was derived by Yidou-Weng et al. (2025) for the case where the constraint is an attribute $s$ predicted by a fully factorized probabilistic classifier $p(s|x_{1:n})$.
@@
We emphasize an advantage of changing only the HMM encoder. Specifically, we compare with an alternative wherein a neural network predicts, given a context $x_{\leq t}$, fresh parameters of a full HMM. Then, the backward computation of $p(\alpha\mid z_{t},s_{t})$ can no longer be precomputed and must be carried out at each decoding step on the freshly predicted HMM, yielding decoding time that scales quadratically in sequence length. Moreover, predicting fresh HMM parameters for each context induces a memory blow-up with batch size. The resulting time and space complexities for training and inference are summarized in Appendix C. A single HMM forward (or backward) step costs $O(H^{2}+HV)$ for dense transition and emission matrices.
@@ 3.2 Architecture
We now specify our hybrid model architectural design: the HMM decoder and the neural network encoder.
@@ 3.2.1 Structured Sparsity
Proposition 3.1.
For any Markov chain $X_{<t}\to Z_{t}\to X_{\geq t}$, we have $I(X_{<t};X_{\geq t})\leq H(Z_{t})\leq\log|\text{supp}\,(Z_{t})|.$
This bound (proven in Appendix A) holds regardless of the encoder $p(z_{t}|x_{<t})$ or decoder $p(x_{\geq t}|z_{t})$, i.e., it holds for standard HMMs as well as neural HMMs. Although Proposition 3.1 provides a strong motivation for increasing the hidden size $H$ to improve the model’s capacity, the number of parameters in an HMM with dense transition and emission matrices grows quadratically in hidden size. Therefore, we also consider using Monarch matrices (Dao et al., 2022), structured matrices that require polynomially fewer parameters in the hidden size yet can still express complex, high-rank matrices, which have been used for scaling tractable models (Zhang et al., 2025a). Monarch matrices for the transition and emission in an HMM use $O(H^{3/2}+H^{1/2}V)$ parameters, yielding a single-step complexity of $O(H^{3/2}+H^{1/2}V)$.
@@ 3.2.2 Neural Network Architectures
Neural HMM performance depends on the encoder architecture, but downstream use also demands low overhead. We therefore reuse the base autoregressive backbone and add only lightweight modules to predict the HMM prior.
Figure 3 shows three variants. (a) Linear head: we freeze the Transformer backbone and learn a linear map from the last hidden state to the latent prior $p(z_{t}\mid x_{\leq t})$. Since the LM hidden states are already computed during decoding, this adds negligible cost. (b) Extra layer: we keep the backbone frozen and insert a small learnable Transformer layer before the linear head, increasing encoder capacity with modest overhead. (c) Finetune: we finetune the transformer backbone jointly with the linear head, maximizing flexibility at the cost of training and inference overhead.
Figure 4: Perplexity of neural HMMs and baseline HMM for varying hidden sizes with dense transition and emission matrices on the left, Monarch matrices on the right.
@@ 4 Experiments
We first assess LTLA as a tractable surrogate by how well it matches the base LM on continuation likelihoods. We then evaluate controlled generation under (i) syntactic constraints (keyword/length DFAs) and (ii) semantic constraints (detoxification via toxicity lookahead). LTLA distills a neural HMM from a fixed base generator (GPT-2 Large for text; Qwen2-VL-2B-Instruct for captioning).
Baselines. For surrogate quality, we compare to a standard HMM and the approach of Lee and Berg-Kirkpatrick (2025). For controlled generation, we compare to prior decoding baselines for syntactic constraints and text detoxification, and to vanilla captioning and prompting for multimodal detoxification; a standard HMM controller isolates the effect of neural prefix encoding.
Reproducibility. Complete training and decoding configurations, including task-specific data construction, hyperparameters, prompts, and qualitative examples, are provided in App. E (GPT-2: Sec. E.1; Qwen2-VL: Sec. E.2.
Research questions. (RQ1) Does neural prefix encoding improve surrogate quality relative to a standard HMM? (RQ2) How does encoder capacity affect this improvement? (RQ3) Do better surrogates yield better controlled generation under syntactic and semantic constraints? (RQ4) Can LTLA condition effectively on multimodal context?
Table 2: CommonGen. HMM-based methods guarantee 100% constraint satisfaction, and LTLA improves generation quality (BLEU-4, ROUGE-L, and CIDEr) over a standard HMM.
Method BLEU-4 ROUGE-L CIDEr Constraint
FUDGE 0.246 0.404 47.0%
A*esque 0.282 0.434 1.52 98.8%
Standard HMM 0.290 0.438 1.55 100.0%
LTLA 0.321 0.454 1.68 100.0%
@@ 4.1 Distillation Performance & Analysis
Figure 4 reports the scaling behavior of neural HMMs compared to standard HMM baselines. The a-axis refers to the hidden size of the HMM. Across both dense and Monarch parameterizations (as introduced in Section 3.2.1), augmenting HMMs with neural components consistently reduces perplexity relative to the baseline. Across the three neural architectures, adding a trainable Transformer layer (i.e., Figure 3(b)) or finetuning GPT (i.e., Figure 3(c)) achieves stronger performance compared to the learnable linear layer. These results demonstrate that neural HMMs scale favorably with hidden dimension, leading to better conditional log-likelihood per token/perplexity. We further observed that the neural encoder achieves better perplexity particularly for shorter continuation lengths, which is to be expected as the first few tokens have the strongest dependence on the context and are where the neural encoder has the most direct effect; see Figure 5 in Appendix D for a plot showing perplexity for varying continuation lengths.
@@
For another reference point, we evaluated the optimization baseline of Lee and Berg-Kirkpatrick (2025), which combines the implicit parameterization of Chiu and Rush (2020) with latent-variable distillation (LVD) (Liu et al., 2023); in Fig. 4 (left) this corresponds to the Implicit-Param HMM curve/bar. In our distillation setting, it does not outperform a carefully tuned standard HMM, likely reflecting the sensitivity of standard HMM training to AdamW hyperparameters, and it remains well behind our neural HMMs.
Table 3: Detoxification for text-only GPT-2 (RealToxicityPrompts) and Qwen2-VL (Hateful Memes).33footnotemark: 3 marks methods with unusually low PPL that also exhibit low diversity, consistent with mode-collapsed “safe” continuations; see §4.2. Full baseline tables are in App. G.
Setting Method Max tox ($\downarrow$) Avg tox ($\downarrow$) dist-2 ($\uparrow$) PPL ($\downarrow$)
GPT-2 text GPT2 baseline 0.385 0.254 0.87 25.57
GeDi 0.363 0.217 0.84 60.03
FUDGE 0.302 0.371 0.78 12.97
PPLM 0.520 0.518 0.86 32.58
DPO 0.180 0.026 0.76 21.59
Standard HMM 0.163 0.016 0.85 29.83
LTLA 0.152 0.015 0.85 29.81
Qwen2-VL Vanilla 0.087 0.249 0.45 2.36
Prompting 0.078 0.224 0.46 2.46
LTLA 0.064 0.188 0.48 3.56
Table 4: Inference-time overhead on GPT-2 detoxification.
Method Ratio
Base LM 1.00
Prompting $\sim$3.0
GeDi / DExperts 2.0–3.0
Mix and Match 7.5
MuCoLa 15–20
PPLM 40.0
LTLA 1.14
@@ 4.2 Controlled Generation
Our neural HMMs both predict future behavior more accurately, which improves benchmark performance, and plug in easily to other models such as VLMs where many existing control methods (including standalone HMMs) are difficult to use. We demonstrate versatility across two constraint families, hard logical constraints and soft semantic attributes, using established constrained-decoding frameworks (Zhang et al., 2024; Yidou-Weng et al., 2025) for both LLMs and VLMs (details in App. E). By learning to look ahead more accurately, we will show that LTLA improves constraint satisfaction, fluency and quality.
@@
Logical constraints. We enforce hard constraints (e.g., keyword constraints) by combining the tractable lookahead model with deterministic finite automata (DFAs) (Zhang et al., 2024). We distill both standard HMMs and LTLA variants from a GPT-2 model finetuned on CommonGen, construct DFAs that accept only sequences satisfying the constraints, then run constrained beam search and select the top hypothesis by base-LM log-likelihood. We evaluate using BLEU-4, CIDEr, ROUGE-L, and perplexity. As shown in Table 2, HMM-based methods are the only ones that guarantee constraint satisfaction (100% versus 47.0–98.8% for prior controllable baselines such as FUDGE and A*esque). Beyond satisfaction, LTLA improves BLEU/ROUGE/CIDEr over a standard HMM and substantially reduces average and maximum perplexity (See App 7, 8), suggesting that the neural prior provides more context-aware guidance under hard constraints without forcing unnatural sequences.
Semantic constraints in LLMs and VLMs.
@@
For semantic control, we couple the tractable lookahead model with a log-linear toxicity classifier trained on RealToxicityPrompts and scored with the Perspective API, following Yidou-Weng et al. (2025). We distill LTLA from the base model (GPT-2 Large for text; Qwen2-VL-Instruct-2B for images). During decoding, the surrogate’s predicted future toxicity provides a lookahead cost that reweights the next-token distribution.
On text-only detoxification, we evaluate GPT-2 Large on RealToxicityPrompts and compare against representative controllable-generation baselines, including trained guides (GeDi, FUDGE), logit-control (PPLM), RL-style approaches (DPO), and the HMM-based controller. On multimodal detoxification, we apply the same mechanism to Qwen2-VL captioning on Hateful Memes, comparing against vanilla captioning and prompt engineering.
Table 3 summarizes the main results (full baseline tables are in App. G). On GPT-2, LTLA reduces maximum toxicity from 0.254 to 0.015 while keeping fluency and diversity comparable to the base model. Some RL-style or guide-based methods report very low perplexity, sometimes even lower than the base model, but this often coincides with sharply reduced diversity, suggesting “overly fluent” mode-collapsed continuations rather than genuinely higher-quality generations (Holtzman et al., 2020). Relative to the standard HMM controller, LTLA further improves toxicity while preserving a similar fluency/diversity profile.
On Qwen2-VL, LTLA reduces maximum toxicity from 0.249 to 0.188 and improves diversity, albeit with higher perplexity. We additionally report long-sequence detoxification in App. G (1k-token generations with 128-step lookahead), where LTLA continues to reduce toxicity while keeping fluency metrics nearly stable, indicating that the benefits extend to longer contexts.
[…]
@@ 5 Related Work and Conclusion
Future-dependent conditional queries generalize next-token prediction and make the LM continuation distribution the key object of interest (Boyd et al., 2022). Since exact conditioning is typically intractable, prior work approximates it via amortized inference, energy-based decoding-time inference, or sequential Monte Carlo over continuations (Hu et al., 2023; Qin et al., 2022; Lew et al., 2023; Zhao et al., 2024; Loula et al., 2025; Ahmed et al., 2025). In parallel, tractable probabilistic models aim to make such conditional inference exact by construction and are actively being scaled in practice (Choi et al., 2020; Liu et al., 2024; Chiu and Rush, 2020; Lee and Berg-Kirkpatrick, 2025). Our contribution is architectural: we learn a globally trained HMM lookahead surrogate and condition only its latent prior on frozen transformer embeddings, aligning with hybrid neural–tractable modeling. (Shao et al., 2020; Dos Martires, 2024).
We proposed Learning to Look Ahead (LTLA), a method for learning tractable continuation distributions that support efficient future-dependent queries while remaining compatible with modern transformer representations. By separating lookback (LM embeddings) from lookahead (a tractable decoder) and learning a prefix-conditioned latent prior, LTLA improves surrogate quality and transfers these gains to downstream control under syntactic and semantic constraints. More broadly, LTLA strengthens the case that continuation modeling can be both accurate and tractable, turning the LM lookahead distribution into an explicit object we can learn, query, and use for probabilistic reasoning.
6 Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.